OpenStudy (0_0youscareme):

translate the following sentence into a equation. A number x subtracted from 36 is three times the sum of four and x

OpenStudy (0_0youscareme):

smh

OpenStudy (anonymous):

A number x subtracted from 36 : 36-x

OpenStudy (anonymous):

"is" - \(=\)

OpenStudy (anonymous):

"three times" - 3* "the sum of four and x" - \((4+x)\)

OpenStudy (anonymous):

I already translated the parts for you. Just make them into one equation. :)

OpenStudy (0_0youscareme):

36-x=3(4+x)?

OpenStudy (anonymous):

Yes.

OpenStudy (0_0youscareme):

tysm for ur help

OpenStudy (anonymous):

No problem! :)

OpenStudy (mathstudent55):

\(\sf \color{red}{A ~number ~x ~subtracted ~from ~36} ~\color{green}{is} ~\color{blue}{three ~times} ~\color{purple}{the ~sum ~of ~four ~and ~x}.\) \(~~~~~~~~~~~~~~~~~~~~~~~\color{red}{36 - x}~~~~~~~~~~~~~~~~\color{green}{=}~~~~~~~~\color{blue}{3 \times }~~~~~~~~~~~\color{purple}{~~~~~~(4 + x)}~~\) Removing the extra spaces and colors: \(36 - x = 3(4 + x)\) You are correct.